OpenStudy (fools101):

What is the remainder when y2 - 18y + 14 is divided by (y - 1)? A) 33 B) 0 C) -3 D) -5

OpenStudy (fools101):

please help!

jimthompson5910 (jim_thompson5910):

just like the last problem in this case, r = 1 now so f(r) = f(1) = 1^2 - 18(1)+14 = 1-18 + 14 =-17 + 14 =-3 So the remainder is -3, choice C

OpenStudy (fools101):

OHHHHH THANK U JIM !!!! : )

jimthompson5910 (jim_thompson5910):

yw